OpenStudy (anonymous):

How Find the domain of the function??? please help........ a. f(x)= 2x, -1 ≤x≤5 b. f(x)= 7/(2x+2) c. g(x)= (x+10)/(x^2-4) d. h(x)= √(10&x^2-6x) e. H(x)= (x^2- 25)/(x+5)

OpenStudy (anonymous):

if the value of function becomes 0/0 , 1/0 , infinity/infinity etc for any x then that x valu will not be included in the domain. eg:- b. f(x)=7/(2x+2) if x = -1 then fx) = 7/0 so the domain would be all real numbers except minus one

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

how about h(x)= √(10&x^2-6x)?

OpenStudy (anonymous):

the value inside √ couldnt be <0 .

OpenStudy (anonymous):

so what is it exactly?

OpenStudy (mathstudent55):

The domain of a function is the the set of values that can be used for the x variable.

OpenStudy (mathstudent55):

A polynomial function has no restrictions on x, so for any polynomial function, the domain is all real numbers.

OpenStudy (anonymous):

can u show the solutions? pls.

OpenStudy (mathstudent55):

A polynomial function is y equals a polynomial in x. Eamples of ploynomial functions: \(y = x + 3\) \(y = 4x\) \(y = x^4 - 8x^3 + 2x^2 - 6x + 7\)

OpenStudy (anonymous):

can u show the solution for this one? h(x)=√(10&x^2-6x)

OpenStudy (mathstudent55):

Where you need to look carefully to find the domain is when the function has a square root, since the radicand can't be negative. That means that the values of x that make the radicand negative have to be excluded from the domain. Also, if there is a denominator in the function, any value of x that will make the denominaotr negative also has to be excluded.

OpenStudy (mathstudent55):

With the above in mind, let's go through your problems.

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

a. f(x) = 2x This is a polynmomial function. There is no value of x that will cause a problem, so the domain is all real numbers.

OpenStudy (anonymous):

get it.

OpenStudy (mathstudent55):

b. f(x)= 7/(2x+2) Here you have a denominator. The problem is that a denominator can't equal zero. All values of x are in the domain except any value of x that will make the denominator zero. To find out which values of x have to be excluded, we set the denominator equal to zero and solve for x. 2x + 2 = 0 2x = -2 x = -1 If x = -1, then 2x + 2 = 2(-1) + 2 = -2 + 2 = 0 Since x = -1 causes the denominator to be zero, then that value of x has to be excluded from the domain. The domain is all real numbers ecept for x = -1.

OpenStudy (mathstudent55):

c. g(x)= (x+10)/(x^2-4) Here we have a denominator again. Once again we set the denominator equal to zero and solve for x.

OpenStudy (mathstudent55):

\(x^2 - 4 = 0\) To solve for x, we factor the left side: \( (x + 2)(x - 2) =0 \) \(x + 2 = 0\) or \(x - 2 = 0\) \(x = -2\) or \(x = 2\) Since -2 and 2 are the values of x that make a zero denominator, they can't be in the domain. The domain is all real numbers except -2 and 2.

OpenStudy (anonymous):

so for letter a, the domain is just all real numbers?

OpenStudy (mathstudent55):

Yes

OpenStudy (anonymous):

if x-2 g(x)=undefined

OpenStudy (anonymous):

is there any solutions for letter a?

OpenStudy (anonymous):

basically the numbers that makes the equation undefined is what you cannot have like for g(x)=(x+20)/(x^2-4)

OpenStudy (anonymous):

yep there's no solution there .

OpenStudy (mathstudent55):

d. h(x)= √(10&x^2-6x) In this case, you don't have a denominator, but you have a square root. The square root of a negative number is not real, so we have to see which values of x will make the radicand nonnegative. First, what is the "&" doing in there? Is this supposed to be \(h(x) =\sqrt{10x^2 - 6x}\) ?

OpenStudy (anonymous):

its h(x)= 10√(x^2-6x) actually.

OpenStudy (mathstudent55):

This? \(h(x) =10\sqrt{x^2 - 6x}\)

OpenStudy (anonymous):

so there's no solution for letter a? the answer is just "the domain is all real numbers?

OpenStudy (anonymous):

yes that's it.

OpenStudy (anonymous):

I thought the answer for letter a is that the domain is all real numbers except 3/5.

OpenStudy (mathstudent55):

For a. the answer is "all real numbers"

OpenStudy (anonymous):

ok thanks

OpenStudy (mathstudent55):

Wait. I have a question on a.

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

a. f(x)= 2x, -1 ≤x≤5

OpenStudy (mathstudent55):

For a. it states above: a. f(x)= 2x, -1 ≤x≤5 Is this "-1 ≤ x ≤ 5" part of the problem ? If so then that is the domain.

OpenStudy (anonymous):

yes it is part of the problem.

OpenStudy (anonymous):

so whats the final answer for a?

OpenStudy (mathstudent55):

Ok, then the answer for a. is "the domain is -1 ≤ x ≤ 5"

OpenStudy (anonymous):

can u show me the solution for letter a pls?

OpenStudy (mathstudent55):

Normally, the domain of f(x) = 2x is all real numbers. In this case, you are told that the domain is -1 ≤ x ≤ 5 when you are given the problem. Answer for a. "the domain is -1 ≤ x ≤ 5"

OpenStudy (anonymous):

alright, thanks.

OpenStudy (mathstudent55):

b. amd c. are already answered above.

OpenStudy (anonymous):

so there is no solutions anymore for letter a?

OpenStudy (mathstudent55):

No, that is it: -1 ≤ x ≤ 5

OpenStudy (anonymous):

ok. so whats the domain for letter d and e?

OpenStudy (mathstudent55):

For d., you need to see for which values of x the radicand is nonnegative. That is the domain. We need to solve the inequality \(x^2 - 6x \ge 0\)

OpenStudy (anonymous):

yeah.

OpenStudy (mathstudent55):

To solve this inequality, first look at its corresponding equation: \(x^2 - 6x = 0\) We solve it by factoring the left side: \(x(x - 6) =0 \) x = 0 or x - 6 = 0 x = 0 or x = 6 The solution to the corresponding equation is x = 0 or x = 6. We need the solution tot he inequality, \( x^2 - 6x \ge 0\), so we draw a number line and plot 0 and 6.

OpenStudy (mathstudent55):